OpenStudy (anonymous):

Propane has a normal boiling point of -42.0 ∘C and a heat of vaporization (ΔHvap) of 19.04 kJ/mol. What is the vapor pressure of propane at 10.0∘C?

OpenStudy (aaronq):

Use the Clasius-Clayperion equation: \(\sf ln(\dfrac{P_2}{P_1})=\dfrac{-\Delta H_{vap}}{R}(\dfrac{1}{T_2}-\dfrac{1}{T_1})\)